Determine the stopping distance for an automobile with an initial speed of 95km and human reaction time of 1s for an acceleration of a=-4.5m/s2, b(a) = -6.7 m/s2
(3A Homework 1 speed of 95 km/h and human reaction time of 1.0s: (a) for arn tni
3A Homework 1 Problem 2.41 Constants | Periodic Table Determine the stopping distances for an automobile with an initial speed of 95 km/h and human reaction time of 1.0 s (a) for an acceleration a - -45 m/s? .(b) for a- -67 m/s
media%2Ff0e%2Ff0ecc7c6-c0b2-4b80-abe4-fd
Answer #1

a )

given

velociity is V = 95 kmph

or

V = 26.39 m/sec

t = 1 sec

a = - 4.5 m/sec2

xo = V x t

= 26.39 x 1

xo = 26.39 m

using equation to find d = u2 - V2 + 2 a xo / 2 a

= 02 - 26.392 + 2 x ( -4.5 ) x 26.39 / 2 x ( -4.5 )

d = - 26.392 + 2 x ( -4.5 ) x 26.39 / 2 x ( -4.5 )

= - 696.4321 - 237.51 / - 9

d = 103.771 m

b )

given

velociity is V = 95 kmph

or

V = 26.39 m/sec

t = 1 sec

b = - 6.7 m/sec2

xo = V x t

= 26.39 x 1

xo = 26.39 m

using equation to find d = u2 - V2 + 2 b xo / 2 b

= 02 - 26.392 + 2 x ( -6.7 ) x 26.39 / 2 x ( -6.7 )

d = - 26.392 + 2 x ( -6.7 ) x 26.39 / 2 x ( -6.7 )

= - 696.4321 - 353.626 / -13.4

d = 78.362 m